OpenStudy (anonymous):

What is ln(1/2) equal?

OpenStudy (jdoe0001):

use das kalkulattor

OpenStudy (anonymous):

-ln(2)

OpenStudy (anonymous):

How does it equal -ln(2)?

OpenStudy (mathstudent55):

ln 1 - ln 2

OpenStudy (usukidoll):

I would use the subtraction log rule first and then compute \[\ln(\frac{1}{2}) \rightarrow \ln 1 - \ln 2 \rightarrow 0 -\ln 2 \rightarrow -\ln 2 \] ln 1 = 0

OpenStudy (anonymous):

Thanks!

OpenStudy (mathstudent55):